OpenStudy (anonymous):

lim as x approaches 0 of (1/x)((1/sqrt(4+x))-(1/2))

OpenStudy (zarkon):

get a common denominator ... multiply top and bottom by conjugate.

OpenStudy (anonymous):

how do u multiply by conjugate if there is no + or - signs?

OpenStudy (zarkon):

get a common denominator ... THEN... multiply top and bottom by conjugate.

OpenStudy (anonymous):

the common denominator i got was (2x^2)(sqrt(4+x)

OpenStudy (zarkon):

\[\lim_{x\to 0}\frac{1}{x}\left(\frac{1}{\sqrt{4+x}}-\frac{1}{2}\right)\] \[\lim_{x\to 0}\frac{1}{x}\left(\frac{2-\sqrt{4+x}}{2\sqrt{4+x}}\right)\] now multiply the top and bottom by \(2+\sqrt{4+x}\).

OpenStudy (anonymous):

oh is that from the top?